# Barron Space Representations for Elliptic PDEs with Homogeneous Boundary Conditions

### v1

@@ 1 Introduction
High-dimensional partial differential equations (PDEs) arise in a wide range of applications, including physics, finance, and control. Classical numerical schemes, such as finite difference and finite element methods, become computationally intractable as the spatial dimension $d$ increases, due to the so-called curse of dimensionality (CoD): to achieve an accuracy of $\varepsilon$, the computational cost scales as $\mathcal{O}(\varepsilon^{-d})$.
@@
Despite their differences in formulation, both DRM and PINNs share the common goal of approximating PDE solutions using neural networks. In practice, the performance of these methods depends critically on several theoretical aspects, including the representational capacity of the chosen neural network architecture, the design and convergence of the training algorithm, and the generalization ability of the resulting model to unseen inputs [SDK20, LLW21, LL22, DRM24]. These considerations naturally give rise to three fundamental theoretical questions in deep learning-based methods:
  • (Approximation) What is the minimal architectural complexity (e.g., depth, width) of a neural network required to approximate a target function with a prescribed accuracy?
  • (Optimization) Given such a network architecture, how can we design training algorithms with provable convergence guarantees to an optimal or near-optimal solution?
  • (Generalization) Under what conditions does the trained network exhibit strong generalization performance, ensuring low error on unseen data?
In this work, we focus on the first question. A prominent framework for addressing this problem is based on Barron functions. In his seminal work [Bar93], Barron proved that any real-valued function on $\mathbb{R}^{d}$ satisfying
$\int_{\mathbb{R}^{d}}|\xi||\widehat{f}(\xi)|{\,\mathrm{d}}\xi<\infty,$ (1.1)
a condition now referred to as having finite Barron norm, can be approximated in the $L^{2}$-norm by a two-layer neural network of the form
@@
with an approximation error independent of the input dimension $d$. This result provides a rigorous theoretical foundation for the ability of neural networks to circumvent the CoD in high-dimensional function approximation.
Following [Bar93], the approximation of Barron functions on bounded domains, such as hypercubes, has been investigated in [KB18, EY18, EMW22, EW22a], where similar dimension-independent approximation results were established. These works collectively suggest that Barron functions, whether defined on the whole space or on bounded domains like hypercubes, can be effectively approximated by two-layer neural networks without suffering from CoD; that is, the required number of neurons does not grow exponentially with the input dimension. This observation inspires a new paradigm for the study of high-dimensional PDEs:
@@
This perspective was first realized by Chen et al. [CLL21], who studied second-order elliptic PDEs on $\mathbb{R}^{d}$. For bounded domains, Marwah et al. [MLLR23] established related results for a broad class of PDEs on the unit hypercube with Dirichlet boundary conditions, under certain regularity assumptions on the coefficients. The approach in [CLL21], motivated by [MLR21], constructs iterative schemes that converge exponentially fast to the solution while maintaining control of the Barron norm through careful analysis at each iteration. A similar methodology was also employed in [MLLR23].
For second-order elliptic PDEs on $\mathbb{R}^{d}$, the Barron function-based approximation framework is essentially complete [CLL21]. On bounded domains, [MLLR23] established results for Dirichlet problems on the unit hypercube, but these require additional assumptions on the Barron coefficients (cf. Remark 2.15). In contrast, the case of second-order elliptic PDEs with Neumann boundary conditions on bounded domains remains largely unaddressed.
@@ 1.1 Our Contributions
    $\begin{cases}-\nabla\cdot A(x)\nabla u+c(x)u=f(x),&\text{in }\Omega,\\ u=0,&\text{on }\partial\Omega,\end{cases}$ (1.3)
@@
    $\begin{cases}-\nabla\cdot A(x)\nabla u+c(x)u=f(x),&\text{in }\Omega,\\ A(x)\nabla u\cdot\nu=0,&\text{on }\partial\Omega,\end{cases}$ (1.4)
where $A(x)$ is a real matrix-valued function, and $\nu$ denotes the outward unit normal vector on $\partial\Omega$. Our main results are stated informally below; precise formulations are given in Theorems 2.13 and 2.17.
@@
using at most $\mathcal{O}(d^{\,C|\log\varepsilon|})$ neurons, where $C>0$ depends only on the coefficients of the PDE.
@@
    $u_{t+1}=u_{t}-\alpha D\mathcal{E}(u_{t}).$
    We show that this scheme converges exponentially fast to the solution $u^{*}$ in the $H^{1}(\Omega)$-norm:
    $\|u_{t}-u^{*}\|_{H^{1}(\Omega)}\lesssim\beta^{t}\quad\text{for all }t\geq 0,$
    for some $0<\beta<1$ depending only on the PDE coefficients. (see Equations (3.19) and (3.20)). In fact, the iterative schemes previously proposed for variational PDEs in [CLL21, MLLR23] can also be interpreted as special cases of the Sobolev gradient flow.
@@
    $\left\|\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}^{2}(\Omega)}}{\sqrt{k}},$
    For the $\operatorname{ReLU}$ activation case, we generalize the strategies in [LLW21] (see Theorem 3.15) and obtain
    $\left\|c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\lesssim\frac{\|g\|_{{\mathcal{B}}^{2}(\Omega)}}{\sqrt{k}}.$
@@ 1.2 Related Works
If the coefficients of the PDE are sufficiently regular such that the solution itself belongs to the Barron class, then it can be approximated by a two-layer neural network without incurring the CoD.
@@ 1.3 Organization
This paper is structured as follows. In Section 2, we formulate the main result. Section 2.1 is devoted to establishing the existence and uniqueness of weak solutions to problems (1.3) and (1.4). In Section 2.2, we introduce the Barron function spaces, which play a central role in our analysis. Building upon these foundations, we present main theorems in Section 2.3. Its proof is given in Section 3, where we provide a detailed exposition of the three-step strategy, with each step elaborated in the corresponding subsection, as briefly outlined in Section 1.1.
2 Formulation and Main Results
In this section, we formulate the problem under consideration and present our main results. Section 2.1 specifies the assumptions that ensure the existence and uniqueness of the weak solution. In Section 2.2, we introduce the Barron spaces and establish a basic consequence of their definition. Finally, Section 2.3 presents the main results of this work.
Throughout the remainder of this work, we let $\Omega\coloneqq(0,1)^{d}\subseteq\mathbb{R}^{d}$ denote the $d$-dimensional unit hypercube. We abbreviate the coefficient functions $A(x)$, $c(x)$, and $f(x)$ by $A$, $c$, and $f$, respectively. The associated second-order elliptic operator is defined as
$\mathcal{L}u\coloneqq-\nabla\cdot A\nabla u+cu.$
2.1 Existence and Uniqueness of Weak Solutions
To ensure the existence and uniqueness of weak solutions to equations (1.3) and (1.4), we impose the following assumption on their coefficients:
Assumption 2.1.
@@ 2.1 Existence and Uniqueness of Weak Solutions
    The real matrix-valued function $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ defined on $\Omega$ is symmetric and uniformly elliptic. By uniform ellipticity, we mean that there exists a constant $a_{\min}>0$ such that
@@
Remark 2.2.
It is straightforward to verify that the finiteness of the operator norm of $A(x)$ implies the finiteness of its $L^{\infty}(\Omega)$-norm,
$\|A(x)\|_{L^{\infty}(\Omega)}\coloneqq\sum_{i,j=1}^{d}\|A_{ij}(x)\|_{L^{\infty}(\Omega)},$
and consequently, each entry $A_{ij}(x)\in L^{\infty}(\Omega)$. However, the converse is not true in general. A counterexample is given by $A(x)=\operatorname{diag}\{\delta_{P}(x),\ldots,\delta_{P}(x)\}$, where $\delta_{P}(x)$ is defined on $\Omega$ and equals $\infty$ if $x=P\coloneqq(1/2,\ldots,1/2)$, and 1 otherwise.
@@
$\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla v+cu_{D}^{*}v-fv{\,\mathrm{d}}x=0\quad\text{for all }v\in H_{0}^{1}(\Omega),$ (2.1)
@@
$\int_{\Omega}\nabla u_{N}^{*}\cdot A\nabla v+cu_{N}^{*}v-fv{\,\mathrm{d}}x=0\quad\text{for all }v\in H^{1}(\Omega).$ (2.2)
Assumption 2.1, together with Remark 2.2 and the Lax–Milgram theorem [Eva10, Theorem 6.2.1], ensures the existence and uniqueness of weak solutions $u_{D}^{*}\in H_{0}^{1}(\Omega)$ and $u_{N}^{*}\in H^{1}(\Omega)$ to (1.3) and (1.4), respectively.
Remark 2.3.
Although, as noted in Remark 2.2, the class of matrix-valued functions with finite $L^{\infty}$-norm is strictly larger than that with bounded operator norm, and the subsequent analysis remains theoretically valid under the $L^{\infty}$-norm (cf. Lemma 3.6 and Theorem 3.4), we choose to adopt the operator norm. This choice is primarily motivated by the desire to ensure that the constants appearing in our main results exhibit milder dependence on the spatial dimension $d$. For example, when $A(x)=I_{d}$ is the $d\times d$ identity matrix, we have
$\|A(x)\|_{\operatorname{op}}=1,\quad\|A(x)\|_{L^{\infty}(\Omega)}=d.$
This demonstrates that the operator norm remains dimension-independent, while the $L^{\infty}$-norm grows linearly with $d$, making the former a more stable and dimension-robust choice.
@@ 2.2 Barron Spaces
Every function in $L^{2}(\Omega)$ admits an $L^{2}$-expansion in family (1) with coefficients in $\mathbb{C}$, though the representation may be non-unique (see Example A.1). Families (2)-(4) form orthogonal bases of $L^{2}(\Omega)$ with coefficients in $\mathbb{R}$.
Remark 2.5.
By an $L^{2}$-expansion of a function in $L^{2}(\Omega)$ (e.g., with respect to family (2)), we mean that there exists a sequence $\{a_{k}\in\mathbb{R}:k\in\mathbb{N}^{d}\}$ such that
$\lim_{N\to\infty}\left\|g(x)-\sum_{\|k\|\leq N}a_{k}S_{k}(x)\right\|_{L^{2}(\Omega)}=0,$
denoted by
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)\quad\text{in }L^{2}(\Omega).$
A collection forms a basis of $L^{2}(\Omega)$ if every $g\in L^{2}(\Omega)$ admits a unique expansion in this family.
Definition 2.6.
Suppose $g(x)$ is a real-valued function defined on $\Omega$ and belongs to $L^{2}(\Omega)$.
@@
    $\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\coloneqq\inf\left\{\sum_{\omega\in\mathbb{Z}^{d}}|a_{\omega}|\left(1+\pi^{n}\|\omega\|^{n}\right):g(x)=\sum_{\omega\in\mathbb{Z}^{d}}a_{\omega}e^{i\pi\omega^{\top}x}\;\text{in }L^{2}(\Omega)\right\}$
@@
Remark 2.7.
The original definition of the Barron norm for functions on $\mathbb{R}^{d}$ was introduced in [Bar93], as given in equation (1.1). In contemporary literature, the version of the Barron norm defined via the Fourier transform—used in, e.g., [SX22, MLLR23, FL25]—is commonly referred to as the spectral Barron norm. In contrast, definitions that avoid the use of the Fourier transform, such as those based on probability measures (e.g., [CLL21, EMW22, SX23]), are typically referred to simply as the Barron norm. Our formulation is inspired by [LLW21, LL22, MLLR23], where cosine basis functions were used in [LLW21, LL22] and exponential basis functions in [MLLR23]. In this work, we introduce the sine-based Barron norm and the $(i,j)$-mixed Barron norm. The sine basis arises naturally when considering functions that vanish on the boundary, while the mixed norm is introduced for certain technical reasons (cf. Lemma 3.12). We also note that if the coefficient matrix $A(x)$ is diagonal, then the mixed Barron norm is not needed. This observation already covers many important families of PDEs, such as the stationary reaction–diffusion equations and static Schrödinger equations.
Remark 2.8.
Once a function $g$ is identified as a Barron function, its expansion in terms of the basis functions from Theorem 2.4 agrees with $g$ almost everywhere. (For the exponential basis, this holds if the weighted sum is suitably chosen to converge.) Consider the sine basis as an example. Suppose $g\in\mathcal{B}_{s}^{n}(\Omega)$, with its $L^{2}$-expansion given by $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$. Then, by the estimate
$\sum_{k\in\mathbb{N}^{d}}|a_{k}S_{k}(x)|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}<\infty,$ (2.3)
the series $\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$ converges absolutely and uniformly on $\Omega$ by the Weierstrass $M$-test, and thus defines a continuous function on $\Omega$. Consequently, it coincides with $g(x)$ almost everywhere on $\Omega$. For simplicity, we will henceforth write
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$
without further qualification.
Remark 2.9.
Once a function $g$ is identified as a Barron function, its Barron norm can be computed using any reordering of its basis expansion. To illustrate, consider the sine basis. Suppose $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$ is an $s$-Barron function of weight $n$, and let $\tau$ be any permutation of $\mathbb{N}^{d}$. Assume further that $g(x)$ admits the expansion
$g(x)=\sum_{k\in\mathbb{N}^{d}}b_{k}S_{\tau(k)}(x)\quad\text{in }L^{2}(\Omega).$
Then, by the uniqueness of the $L^{2}(\Omega)$ expansion, we have $b_{k}=a_{\tau(k)}$ for all $k\in\mathbb{N}^{d}$, and hence
$\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}=\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|a_{\tau(k)}|\left(1+\pi^{n}\|\tau(k)\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|b_{k}|\left(1+\pi^{n}\|\tau(k)\|^{n}\right),$
where the second equality is justified by the absolute convergence of the series. This reordering invariance will be used in the sequel, including in the proofs of Proposition 3.10 and Lemma 3.12.
@@
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}\frac{a_{k}}{(2i)^{d}}\prod_{j=1}^{d}\left(e^{i\pi k_{j}x_{j}}-e^{-i\pi k_{j}x_{j}}\right)=\sum_{k\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\pm\frac{a_{k}}{(2i)^{d}}e^{i\pi v_{m}^{\top}x},$
where each vector $v_{m}=(\pm k_{1},\dots,\pm k_{d})\in\mathbb{Z}^{d}$ corresponds to one of the $2^{d}$ sign combinations. By the reordering invariance of the Barron norm (Remark 2.9) and the definition of the $e$-Barron norm, we obtain
$\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\left|\frac{a_{k}}{(2i)^{d}}\right|\left(1+\pi^{n}\|v_{m}\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)},$
where we used that $\|v_{m}\|=\|k\|$.
@@
In particular, the following inclusions hold for all $n\geq 0$:

### v2

@@ 1 Introduction
High-dimensional partial differential equations (PDEs) arise in a wide range of applications, including physics, finance, and control. Classical numerical schemes, such as finite difference and finite element methods, become computationally intractable as the spatial dimension $d$ increases, due to the so-called curse of dimensionality (CoD): to achieve an accuracy of $\varepsilon$, the computational cost scales as $\mathcal{O}(\varepsilon^{-d})$, exponentially in $d$.
@@
Despite differences in formulation, both DRM and PINNs share the goal of learning PDE solutions with neural networks. Their practical performance hinges on several theoretical factors: the representational capacity of the network architecture, the design and convergence of the training algorithm, and the generalization ability of the trained model to unseen inputs [SDK20, LLW21, LL22, DRM24]. These considerations motivate three fundamental theoretical questions for deep learning–based PDE methods:
  • Approximation: What is the minimal architectural complexity (e.g., depth, width) required to approximate a target function to a prescribed accuracy?
  • Optimization: For a given architecture, how can we design training algorithms with provable convergence to an optimal or near-optimal solution?
  • Generalization: Under what conditions does the trained network generalize effectively, achieving low error on unseen data?
In this work, we focus on the first question. A prominent framework for addressing this problem is based on Barron functions. In his seminal work [Bar93], Barron proved that for any real-valued function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, if its Fourier transform $\widehat{f}$ satisfies
$\int_{\mathbb{R}^{d}}\|\xi\|\cdot|\widehat{f}(\xi)|{\,\mathrm{d}}\xi<\infty,$ (1.1)
a condition now referred to as having finite Barron norm, then $f$ can be approximated in the $L^{2}$-norm by a two-layer neural network of the form
@@
with an approximation error independent of the input dimension $d$. This result provides a rigorous theoretical foundation for the ability of neural networks to circumvent the CoD in high-dimensional function approximation; that is, the required number of neurons does not grow exponentially with the input dimension. Following [Bar93], there have been multiple variants and generalizations of Barron functions with dimension-independent approximation results of a similar spirit, see e.g. [KB18, EMW22, EW22a, SX22, SX23, SX24].
This observation inspires a new paradigm for the study of high-dimensional PDEs:
@@
This perspective was first realized by Chen et al. [CLL21], who developed a Barron function-based approximation framework for second-order elliptic PDEs on the whole space $\mathbb{R}^{d}$. Their iterative schemes, motivated by [MLR21], converge exponentially fast while maintaining control of the Barron norm. On bounded domains, Marwah et al. [MLLR23] extended this methodology to Dirichlet problems on the unit hypercube, though their results require additional regularity assumptions on the coefficients (see the discussion of Theorem 2.7 in Section 2.2). In contrast, the case of second-order elliptic PDEs with Neumann boundary conditions on bounded domains remains largely open.
@@ 1.1 Our Contributions
[…]
@@
[…]
where $A(x)$ is a real matrix-valued function, and $\nu$ denotes the outward unit normal vector on $\partial\Omega$. Our main results are stated informally below; precise formulations are given in Theorems 2.7 and 2.9.
@@
using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where $C>0$ depends only on the Barron norms of the PDE coefficients.
@@
    $u_{t+1}=u_{t}-\alpha D\mathcal{E}(u_{t}),$
    where $\mathcal{E}$ is the energy functional associated with (1.3) or (1.4). We show that this scheme converges exponentially fast to the solution $u^{*}$ in the $H^{1}(\Omega)$-norm:
    $\|u_{t}-u^{*}\|_{H^{1}(\Omega)}\lesssim\beta^{t},\quad\text{for all }t\geq 0,$
    for some $0<\beta<1$ depending only on the PDE coefficients (see Equations (3.17) and (3.18)). In fact, the iterative schemes previously proposed for variational PDEs in [CLL21, MLLR23] can also be interpreted as special cases of the Sobolev gradient flow.
@@
    $\left\|\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}(\Omega)}}{\sqrt{k}},$
    where $\|g\|_{\mathcal{B}(\Omega)}$ is the Barron norm of $g$. For the $\operatorname{ReLU}$ activation case, we generalize the strategies in [LLW21] (see Theorem 3.11) and obtain
    $\left\|c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\lesssim\frac{\|g\|_{{\mathcal{B}}(\Omega)}}{\sqrt{k}}.$
@@ 1.2 Related Works
If the coefficients of the PDE are sufficiently regular in the Barron sense, does the solution belong to the Barron class?
@@ 1.3 Organization
The rest of this paper is organized as follows. Section 2 present our main results after introducing the Barron spaces and basic properties. The proofs, based on a three-step strategy outlined in Section 1.1, are detailed in Section 3.
2 Main Results
In this section, we present our main results. To ensure the existence and uniqueness of weak solutions to PDEs (1.3) and (1.4), we impose the following minimal assumption on their coefficients. All functions in this paper are real-valued unless otherwise noted.
Assumption 2.1 (Coefficients of Elliptic PDEs).
@@ 2 Main Results
    The matrix-valued function $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ defined on $\Omega$ is symmetric and uniformly elliptic. By uniform ellipticity, we mean that there exists a constant $a_{\min}>0$ such that
@@
$\int_{\Omega}\left(\nabla u_{D}^{*}\cdot A\nabla v+cu_{D}^{*}v-fv\right){\,\mathrm{d}}x=0\quad\text{for all }v\in H_{0}^{1}(\Omega),$ (2.1)
@@
$\int_{\Omega}\left(\nabla u_{N}^{*}\cdot A\nabla v+cu_{N}^{*}v-fv\right){\,\mathrm{d}}x=0\quad\text{for all }v\in H^{1}(\Omega).$ (2.2)
Assumption 2.1 and the Lax–Milgram theorem [Eva10, Theorem 6.2.1] ensure the existence and uniqueness of weak solutions $u_{D}^{*}\in H_{0}^{1}(\Omega)$ and $u_{N}^{*}\in H^{1}(\Omega)$ to (1.3) and (1.4), respectively.
@@ 2.1 Barron Spaces
Every function in $L^{2}(\Omega)$ admits an $L^{2}$-expansion in family (1) with coefficients in $\mathbb{C}$, though the representation may be non-unique. Families (2)-(4) form orthogonal bases of $L^{2}(\Omega)$ with coefficients in $\mathbb{R}$.
Definition 2.3 (Barron Norms and Barron Functions).
Suppose $g(x)\in L^{2}(\Omega)$.
@@
    $\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\coloneqq\inf\left\{\sum_{\omega\in\mathbb{Z}^{d}}|a_{\omega}|\left(1+\pi^{n}\|\omega\|^{n}\right):g(x)=\sum_{\omega\in\mathbb{Z}^{d}}a_{\omega}e^{i\pi\omega^{\top}x}\;\text{in }L^{2}(\Omega)\right\}.$
@@
The original definition of the Barron norm for functions on $\mathbb{R}^{d}$ was introduced in [Bar93], as given in equation (1.1). In contemporary literature, the version of the Barron norm defined via the Fourier transform, used in, e.g., [SX22, MLLR23, FL25], is commonly referred to as the spectral Barron norm. In contrast, definitions that avoid the use of the Fourier transform, such as those based on probability measures (e.g., [CLL21, EMW22, SX23]), are typically referred to simply as the Barron norm. Our formulation is inspired by [LLW21, LL22, MLLR23], where cosine basis functions were used in [LLW21, LL22] and exponential basis functions in [MLLR23]. In this work, we introduce the sine-based Barron norm and the $(i,j)$-mixed Barron norm. The sine basis arises naturally when considering functions that vanish on the boundary, while the mixed norm is introduced for certain technical reasons (cf. Lemma 3.8). We also note that if the coefficient matrix $A(x)$ is diagonal, then the mixed Barron norm is not needed. This observation already covers many important families of PDEs, such as the stationary reaction–diffusion equations and static Schrödinger equations.
@@
$g(x)$ $=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}\frac{a_{k}}{(2i)^{d}}\prod_{j=1}^{d}\left(e^{i\pi k_{j}x_{j}}-e^{-i\pi k_{j}x_{j}}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\frac{a_{k}}{(2i)^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\prod_{j=1}^{d}\epsilon_{j}\cdot e^{i\pi(\epsilon\circ k)^{\top}x},$
where $\epsilon\circ k$ is the Hadamard product of $\epsilon$ and $k$. This observation leads to that
$\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|\epsilon\circ k\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}.$
@@
In particular, the following inclusions hold: